OpenStudy (anonymous):

Let f(x) = Integral cos xdx for 0

OpenStudy (amistre64):

so what is the integration?

OpenStudy (anonymous):

cos xdx?

OpenStudy (amistre64):

thats the arguement. f(x) equals the integral of cos x dx. What does the integration give us?

OpenStudy (anonymous):

Total area under the curve?

OpenStudy (amistre64):

still trying to work out the rest of what you are asking really; can you provide a better representation of the question? ascii doesnt do well with it

OpenStudy (amistre64):

sin(x) + C derives to cos(x) so we have an antiderivative to play with; but the information in the question is not making a lot of sense to me at the moment wihtout some clarification

OpenStudy (anonymous):

\[Let f(x) = ∫\cos x dx for 0 < x < \pi. If f ( \pi/6) = 1, then f ( \pi / 2) ?\]

OpenStudy (anonymous):

That's exactly what the question stated.

OpenStudy (amistre64):

if you can take a screenshot or picture of the question and attach it to your post, it would be easy to see if you are presenting the information correctly to us. what im used to seeing is something like: \[f(x)=\int_{0}^{h(x)}g(t)~dt\] or some such ....

OpenStudy (anonymous):

The question that I stated above. is correct. There isn't any information given other the that..

OpenStudy (anonymous):

I think I will have take 2 integrals. One from pi/6 to pi and then another one from pi/2 to pi

OpenStudy (amistre64):

otherwise, all im reading this as is: \[f(x)=\int_{0}^{\pi}~cos(x)~dx=0\] therefore f(x) = 0 for any x between 0 and pi ....

OpenStudy (anonymous):

\[\int\limits_{\pi/6}^{\pi} - \int\limits_{\pi}^{\pi/2} \]

OpenStudy (anonymous):

And then then take the sin of those variables and then add them together? Would it work?

OpenStudy (amistre64):

if its an intial value problem such that the antiderivative of cos(x) = sin(x)+C; then sin(pi/6)+C=1 sin(pi/2) - sin(pi/6) = ? youre new rendition is even stranger ... which is why im thinking this might be simpler if you could provide a "hardcopy" of the question to make sure.

OpenStudy (anonymous):

http://www.mrefram.com/apcalc/APCalcReleasedFinal.pdf. Gp that website and look at number 8.

OpenStudy (anonymous):

the you will find the question

OpenStudy (amistre64):

its saying its a bad link ....

OpenStudy (anonymous):

http://www.mrefram.com/apcalc/APCalcAssignments.htm

OpenStudy (anonymous):

Then click on practice final review. If possible..

OpenStudy (amistre64):

got it

OpenStudy (anonymous):

Alright :D

OpenStudy (amistre64):

is that they answer key below it?

OpenStudy (anonymous):

Yes, the answer is suppose to be 3/2

OpenStudy (amistre64):

what i do notice is that it is an indefinite integral; no limits. So im thinking my initial value idea is more spot on.

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

\[f(x)=\int~cos(x)~dx\] \[f(x)=sin(x)+C\] \[when~x=pi/6:~sin(pi/6)+C=1\] this allows us to find a suitable value for C; C = 1 - sin(pi/6) therefore:\[f(x)=sin(x)+1-sin(pi/6)\]

OpenStudy (amistre64):

when x=pi/2, what do we get?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I got 3/2 for the answer

OpenStudy (amistre64):

me too; the key was in understanding what type of integral they were asking for. a definite integral with limits from a to b defines the area underneath a curve. an indefinite integral provides us a family of antiderivatives which we can use an initial value on to find a specific function to play with.

OpenStudy (anonymous):

Alright. Thank you so much. I got it :D

OpenStudy (amistre64):

youre welcome :)